# The Connection Between Monetary Policy and Housing Prices: Public Perception and Expert Communication

### v1

@@ 1 Introduction
Fourth, we contribute to the yet very small body of evidence on heterogenous impact of expert commentaries on the public opinion. hirschman2014economists discuss the general impact of economists’ institutional positions on authority and shaping policy, and find that expert economists appear to receive high professional authority. dommett2019we also rejects the ‘post-truth’ argument in an assessment of survey data from the UK and continental Europe as the vast majority appeared to see a value of experts’ knowledge and policy advise. Yet, the main finding from this review is that there is currently too little understanding of the public perception of different types of experts. Such heterogeneity across different experts’ opinions are more rarely studied: One notable exception is the study by sievertsen2025female that detects a gender gap in credibility within the group of academic economists publicly providing expertise and advise. In the present study, we aim to contribute to this strand of literature by studying heterogeneity of the credibility of experts’ opinions across academic and policy economists. We find that participants are more reactive to information provided by academic economists compared to central bankers. This result suggests that the public may be more reluctant to acknowledge expertise by central bankers as compared to academic economists, with important implications for central bank communication.
Fifth, we contribute to the general literature of financial literacy and implications for decision-making. General levels of financial literacy and education attainment have been shown to strongly correlate with the quality of financial decision-making including the likelihood of mortgage stress (hu2024financial), high-cost borrowing (lusardi2015debt) or inadequate financial planning (lusardi2014economic). In this spirit, we assess the association between specifically monetary policy literacy and acquired expertise via past active involvement in housing markets and particularly transactions. Own involvement in real estate transactions including a mortgage are a clear predictor of CMP literacy yet not UMP. This notion is in-line with findings by bucks2008borrowers documenting that mortgage-holders all appear to have a basic understanding of their mortgage terms yet fail in large proportions concerning more complex product features and appear often unprepared for unfavourable future interest rate scenarios.
@@ 4.2.1 Set-up
Communication can shape expectations. Central bank communication indeed influences financial markets (swanson2021measuring) and information about inflation can shape households’ future inflation expectations (coibion2022monetary; hoffmann2021effects). Central bank communication thus aims to make use of this and has accordingly improved substantially over the last decades.
@@
”There are central bankers arguing that when interest rates fall, house prices will rise. For example, Deputy Governor Chen Nan-kuang of Taiwan’s central bank states that “loose monetary policy [note: meaning low interest rates] is indeed one of the main reasons for rising house prices.” (Source: The Taiwan Banker NO.145101010Chen2022_HousingPricesFinancialStability.)
@@ 4.2.3 Does It Matter Who provides Experts Advice?
We start with assessing the importance of the type of economist providing expert advise. 7(d) therefore splits the information effect by information provider: a central banker or an academic economist. We first focus on the group of participants updating their response from an initially wrong to a correct answer (“good switcher”).
@@
Overall, these results suggest that information provided by central bankers is somewhat less effective in aligning perceptions among the general public with economic theory than information provided by a research economist. While further research is needed to confirm the robustness of these results across countries and different agents acting as experts, this result may be a first hint that MP communication may be most effective when relying on a strategic mix of experts including academic economists focusing on theoretical background information.
@@ 5 Conclusions
Our results further reveal important heterogeneity in how expert information is received. Participants respond more strongly to statements by academic economists than to those by central bank economists. This suggests a credibility gap that is highly relevant for monetary authorities. While central banks are the primary communicators of monetary policy, they may not always be perceived as the most trusted source of economic expertise by the general public. This finding calls for a rethinking of communication strategies, including closer cooperation with academic institutions and a stronger emphasis on transparency and independence in messaging.
@@ 6 Declaration of Generative AI and AI-assisted technologies in the Manuscript Preparation Process
Notes: SE characteristics refers to the same set as used in Model (1) in Table 2 and Table 3.

### v2

@@ 1 Introduction
Fourth, we contribute to the yet very small body of evidence on heterogenous impact of expert commentaries on the public opinion. hirschman2014economists discuss the general impact of economists’ institutional positions on authority and shaping policy, and find that expert economists appear to receive high professional authority. dommett2019we also rejects the ‘post-truth’ argument in an assessment of survey data from the UK and continental Europe as the vast majority appeared to see a value of experts’ knowledge and policy advise. Yet, the main finding from this review is that there is currently too little understanding of the public perception of different types of experts. Such heterogeneity across different experts’ opinions are more rarely studied: One notable exception is the study by sievertsen2025female that detects a gender gap in credibility within the group of academic economists publicly providing expertise and advise. In the present study, we aim to contribute to this strand of literature by studying heterogeneity of the credibility of experts’ opinions across academic and policy economists. We find that participants are more reactive to information provided by academic economists compared to central bankers. We do not detect heterogeneities in this effect related to participants’ stated trust in institutions and perceived corruption, democratic participation, education attainment, or other dimensions. This general result therefore suggests that the public may be more reluctant to acknowledge expertise by central bankers as compared to academic economists, with important implications for central bank communication.
Fifth, we contribute to the general literature on financial literacy and its implications for decision-making. General levels of financial literacy and education attainment have been shown to strongly correlate with the quality of financial decision-making including the likelihood of mortgage stress (hu2024financial), high-cost borrowing (lusardi2015debt) or inadequate financial planning (lusardi2014economic). In this spirit, we assess the association between specifically monetary policy literacy and acquired expertise via past active involvement in housing markets and particularly transactions. Own involvement in real estate transactions including a mortgage are a clear predictor of CMP literacy yet not UMP. This notion is in-line with findings by bucks2008borrowers documenting that mortgage-holders all appear to have a basic understanding of their mortgage terms yet fail in large proportions concerning more complex product features and appear often unprepared for unfavourable future interest rate scenarios.
@@ 4.2.1 Set-up
Communication can shape expectations. Central bank communication indeed influences financial markets (swanson2021measuring) and information about inflation can shape households’ future inflation expectations (coibion2022monetary; hoffmann2021effects). Central bank communication thus aims to make use of this and has accordingly been a rapidly growing area of research (see ehrmann2025trust, for a survey of this literature).
@@
“There are central bankers arguing that when interest rates fall, house prices will rise. For example, Deputy Governor Chen Nan-kuang of Taiwan’s central bank states that “loose monetary policy [note: meaning low interest rates] is indeed one of the main reasons for rising house prices.” (Source: The Taiwan Banker NO.145101010Chen2022_HousingPricesFinancialStability.)
@@ 4.2.3 Does It Matter Who provides Experts Advice?
Next, we assess the importance of the type of economist providing expert advise. 7(d) therefore splits the information effect by information provider: a central banker or an academic economist. We first focus on the group of participants updating their response from an initially wrong to a correct answer (“good switcher”).
@@
We further test whether there are heterogeneities in this tendency to follow advise by academic economists more or whether this is a general finding. Concretely, we first make use of stated trust in the government and institutions,141414(S1) How good do you think the government and tax authorities are at things like efficient use of resources, avoidance of mistakes and preventing fraud? perceptions of corruption,151515(S2) Please tell me whether you think the government and tax authorities in your country give special advantages to certain people or deal with everyone equally? and political participation in the form of voting turnout which has been shown to correlate with trust in policy and politics (Hooghe2017TrustAndElections) in general. Further, we assess heterogeneity by education attainment (tertiary vs. less), and expectation or past receipt of an inheritance161616(S3) Do you or any other member of your household expect to receive an inheritance or a substantial gift (worth more than GBP 8,500), including money or any other assets (e.g., houses, land, cars, etc.) in the future?; (S4) Have you or any member of your household ever received an inheritance or a substantial gift, including money or any other major assets (e.g., houses, land, cars, etc.)?.
We split all dimensions – that are not binary in nature – into two categories (high/low) to sustain sufficient observations per group, and interact the binary outcome with the type of information provider. Table 13 reports results for the direct question (Question 4), and Table 14 for the indirect question (Question 5).
For no single dimension, the interaction term is significantly different from zero, no matter whether we include controls or not, and, hence, we conclude that there is no systematic deviation among the dimensions tested from the general result that consumers are more likely to follow academic economists’ advise in this setting.
Overall, these results indicate that information conveyed by central bankers is somewhat less effective at aligning public perceptions with economic theory than information provided by research economists. Moreover, the absence of heterogeneous treatment effects suggests that this pattern holds broadly across different groups, underscoring the generality of the finding.
While further research is needed to assess the robustness of these results across countries, over time, and with different types of expert communicators, our findings provide an initial indication that monetary policy communication may be most effective when it draws on a strategic mix of experts, including academic economists who can emphasize and clarify the underlying theoretical foundations.
@@ 5 Conclusions
Our results further reveal important heterogeneity in how expert information is received. Participants respond more strongly to statements by academic economists than to those by central bank economists. This suggests a credibility gap that is highly relevant for monetary authorities. While central banks are currently the primary communicators of monetary policy, they may not always be perceived as the most trusted source of economic expertise by the general public. This finding calls for a rethinking of communication strategies, including closer cooperation with academic institutions and a stronger emphasis on transparency and independence in messaging.
@@ 6 Declaration of Generative AI and AI-assisted technologies in the Manuscript Preparation Process
Notes: SE refers to the same set as used in Model (1) in Table 2 and Table 3.
Table 13: Heterogeneity Analysis – Direct Effect
Dependent variables: Incorrigibles, Susceptibles, Experts
Trust Corruption Vote
(1) (2) (3) (4) (5) (6)
Academic Economist Info $0.177$ 0.233* 0.2340.234*** 0.232*** 0.481* 0.516**
($0.124$) (0.128) (0.019) (0.087) (0.246) (0.252)
Group $\times$ Academic Economist Info 0.007-0.007 -0.091 -0.206 -0.391* -0.279 -0.310
($0.227$) (0.151) (0.213) (0.223) (0.213) (0.218)
SE
Observations 3,729 3,453 3,729 3,454 3,729 3,454
Log Likelihood -3,315.796 -3,195.920 -3,242.800 -3,153.071 -3,314.906 -3,195.428
AIC 4,893.216 6,419.8 4,893.881 6,334.1 4,892.375 6,416.9
Education Expected inheritance Past inheritance
(1) (2) (3) (4) (5) (6)
Academic Economist Info 0.109 0.249 0.427 0.438 0.355 0.306
(0.201) (0.206) (0.270) (0.206) (0.249) (0.255)
Group $\times$ Academic Economist Info 0.048 -0.058 -0.145 -0.156 -0.107 -0.083
(0.141) (0.144) (0.151) (0.154) (0.143) (0.146)
SE
Observations 3,467 3,454 3,463 3,445 3,463 3,445
Log Likelihood -3,302.202 -3,196.363 -3,299.094 -3,189.868 -3,296.536 -3,189.211
AIC 6,614.4 6,418.7 6,608.2 6,407.7 6,603.1 6,406.4
Notes: Ordered Logit Models on the outcome in Question 4. SE refers to the same set as used in Model (1) in Table 2 and Table 3.
Table 14: Heterogeneity Analysis – Indirect Effect
Dependent variables: Incorrigibles, Susceptibles, Experts
Trust Corruption Vote
(1) (2) (3) (4) (5) (6)
Academic Economist Info $0.132$ 0.137 0.233*** 0.226*** 0.289 0.377
(0.126) (0.082) (0.082) (0.084) (0.241) (0.248)
Group $\times$ Academic Economist Info 0.040 0.060 -0.220 -0.138 -0.114 -0.175
(0.149) (0.153) (0.143) (0.147) (0.207) (0.212)
SE
Observations 3,488 3,464 3,489 3,465 3,488 3,465
Log Likelihood -3,257.630 -3,117.796 -3,257.139 -3,117.709 -3,248.483 -3,117.853
AIC 6,525.3 6,263.6 6,524.3 6,263.4 6,507.0 6,261.7
Education Expected inheritance Past inheritance
(1) (2) (3) (4) (5) (6)
Academic Economist Info -0.065 -0.007 0.171 0.243 -0.040 0.056
(0.203) (0.208) (0.281) (0.288) (0.252) (0.260)
Group $\times$ Academic Economist Info 0.187 0.143 -0.003 -0.037 0.119 0.071
(0.146) (0.149) (0.156) (0.160) (0.144) (0.149)
SE
Observations 3,478 3,465 3,474 3,456 3,474 3,456
Log Likelihood -3,224.417 -3,117.730 -3,233.376 -3,103.513 -3,241.555 -3,109.541
AIC 6,458.8 6,261.5 6,476.8 6,235.0 6,493.1 6,247.1
Notes: Ordered Logit Models on the outcome in Question 5. SE refers to the same set as used in Model (1) in Table 2 and Table 3.